# Evaluation of the absolute single-photon detection efficiency of HRPPD

### v1

@@ 4.1 Statistical and systematic uncertainties
Statistical uncertainties quoted in Table 1 follow from fit errors of the photodiode calibration curves (see Section 3), spread of the reference photodiode current measurements, binomial error when calculating a fraction of events with HRPPD pulses above threshold, as well as its error propagation in formula 3.
We assign a systematic uncertainty of 0.8% to this PDE estimate, reflecting measured quantity fluctuations during a setting up stage when the components were mounted and dismounted a number of times before the final data set was taken, effect of a threshold variation when calculating number of HRPPD pulses, as well as a 1.65% relative uncertainty of the Newport photodiode calibration.
@@ 5 HRPPD QE measurement
We performed the QE measurement using both the PiLas laser (at a variety of tunes and repetition rates) and the monochromator, in order to see a possible difference between a pulsed and a continuous light source. A configuration shown in Figure 1 (right) was used in both cases, without changing anything inside of the dark box. The photocathode bias voltage was 200 V, provided by the Keithley 6487 picoammeter. The results are listed in Table 2. The light source was turned off for about ten minutes between the successive measurements, which were taken in a sequence as listed in the table.
@@
Laser @ 40 MHz tune 0% $\sim$2006 nA 23.1 $\pm$ 0.3 %
Laser @ 30 MHz tune 0% $\sim$1575 nA 22.7 $\pm$ 0.3 %
Laser @ 20 MHz tune 0% $\sim$1173 nA 22.9 $\pm$ 0.3 %
Laser @ 10 MHz tune 0% $\sim$625 nA 23.3 $\pm$ 0.2 %
Laser @ 1 MHz tune 0% $\sim$56 nA 24.3 $\pm$ 0.1 %
Laser @ 40 MHz tune 0% $\sim$1916 nA 22.0 $\pm$ 0.4 %
Laser @ 40 MHz tune 60% $\sim$957 nA 22.7 $\pm$ 0.3 %
Laser @ 40 MHz tune 80% $\sim$634 nA 23.0 $\pm$ 0.2 %
Laser @ 40 MHz tune 93% $\sim$340 nA 23.3 $\pm$ 0.3 %
Monochromator $\sim$63 nA 22.9 $\pm$ 0.1 %
@@
One can see that the measured values show a certain trend in the laser data, namely the QE estimate goes up with decreasing HRPPD photocathode current. However, the monochromator point (continuous wave beam, but also low intensity) is not really consistent with this observation.
We should also mention that a Hamamatsu calibration coefficient came out slightly different for the monochromator case, as compared to the laser light source, which fact we tentatively attribute to a usage of a pellicle beam splitter and a possible small residual polarization of the laser light after coupling into the multi-mode fiber (while the monochromator light is expected to be unpolarized). For the purposes of this report, we are therefore taking a QE value of 24.3 $\pm$ 0.1 (stat) %, obtained with a PiLas laser at the lowest intensity presented in Table 2, with the understanding that this number may have a substantial systematic uncertainty up to at least $\pm$ 0.5 % associated with it.
@@ 6 HRPPD Collection Efficiency
The Collection Efficiency follows from the PDE and QE measurements at the same wavelength and at the same physical location in the HRPPD active area as CE = PDE/QE. Substituting the numbers obtained in the last two sections, one can see that the HRPPD CE is about 79 %. This number looks reasonable given that most of the primary photoelectrons hitting the interstitial space between the pores of the top surface of the first MCP either do not bounce back or do not make it into the pores with an energy sufficient to initiate an avalanche. For the MCPs used in EIC HRPPDs, the pores occupy about 70 % of the surface area, in a reasonable agreement with our CE estimate.
The proposed technique also allows one to use timing information on event per event basis in order to disentangle contributions of ”prompt” photoelectrons and those which bounced off the interstitial space and initiated the avalanche with a delay of up to several hundred picoseconds. The fraction of events in the core out of the total signal events is measured to be approximately 97.5%.

### v2

@@ 5.1 Statistical and systematic uncertainties
Statistical uncertainties quoted in Table 1 follow from the spread of the reference photodiode current measurements and binomial error when calculating a fraction of events that have HRPPD pulses above threshold with its error propagation in formula 3.
The following sources of systematic uncertainty are taken into account: a 1.65% relative uncertainty of the Newport photodiode calibration, and a 0.32% relative uncertainty of the attenuation factor of the OD5 filter from the fit errors of the photodiode calibration curves (see Section 4).
@@ 6 HRPPD QE measurement
We performed the QE measurement using both the PiLas laser (at a variety of repetition rates and tunes) and the monochromator, in order to see a possible difference between a pulsed and a continuous light source. A configuration shown in Figure 1 (right) was used in both cases, without changing anything inside of the dark box. The photocathode bias voltage was 200 V, provided by the Keithley 6487 picoammeter. The results are listed in Table 2. The light source was turned off for about ten minutes between the successive measurements.
@@
Laser @10 MHz tune 0% 3043$\pm$35 nA 22.88$\pm$0.26 %
Laser @40 MHz tune 98% 1069$\pm$13 nA 22.95$\pm$0.46 %
Laser @1 MHz tune 0% 267.47$\pm$0.70 nA 23.38$\pm$0.06 %
Laser @0.5 MHz tune 0% 132.03$\pm$0.18 nA 23.67$\pm$0.04 %
Laser @0.1 MHz tune 0% 26.46$\pm$0.01 nA 23.76$\pm$0.01 %
Laser @50 kHz tune 0% 13.23$\pm$0.02 nA 24.37$\pm$0.05 %
Laser @30 kHz tune 0% 7.99$\pm$0.01 nA 24.37$\pm$0.03 %
Monochromator 231.6$\pm$2.3 nA 24.05$\pm$0.39 %
@@
One can see that the measured values exhibit a certain trend in the laser data, specifically that the QE estimates increase with decreasing HRPPD photocathode currents. We stop at 30 kHz since the photo-current has decreased to a similar magnitude of the leakage current ($\sim$6 nA) and the result remained identical to what we obtained at 50 kHz.
We are therefore taking a QE value of 24.37 $\pm$ 0.03 (stat) %, obtained with a PiLas laser at the lowest intensity presented in Table 2, as final result with a substantial systematic uncertainty 0.32% taken from the difference between the laser measurement and the monochromator measurement.
@@ 7 HRPPD Collection Efficiency
The Collection Efficiency follows from the PDE and QE measurements at the same wavelength and at the same physical location in the HRPPD active area as CE = PDE/QE. Substituting the numbers obtained in the last two sections, one can see that the HRPPD CE is 70.3 $\pm$ 1.4%, as measured at a particular active area spot. This number looks reasonable given that most of the primary photoelectrons hitting the interstitial space (see Figure 11) between the pores of the top surface of the first MCP either do not bounce back or do not make it into the pores with an energy sufficient to initiate an avalanche. For the MCPs used in EIC HRPPD #27, the pores occupy 69.4 % of the surface area [8], in a reasonable agreement with our CE estimate.
Figure 11: An MCP sample image from the same capillary glass array block as used for manufacturing the HRPPD #27, taken by a scanning electron microscope, shows the configuration of the $\sim 10\mu m$ diameter pores and the interstitial space between them.
8 Conclusions
The photon detection efficiency of one HRPPD from the first batch dedicated for EIC has been measured to be 17.13 $\pm$ 0.11 (stat) $\pm$ 0.29 (sys) % at 398.6 nm, in a photoelectron pulse counting mode using a picosecond diode laser. HRPPD QE at the same spot and at the same wavelength has been evaluated to be 24.37 $\pm$ 0.03 (stat) $\pm$ 0.32 (sys) %, leading to a CE estimate of 70.3 $\pm$ 1.6%, which is consistent with the expectations.
Acknowledgements
We are grateful to Dr. Zhoudunming Tu and the Center for Functional Nanomaterials at BNL for their assistance with the scanning electron microscope imaging and to Bob Azmoun for his contribution to the setup commissioning.